Circular No. 7279

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1859+226 R. M. Wagner, Ohio State University; P. S. Smith, National Optical

Astronomy Observatory; G. D. Schmidt, University of Arizona; and C. R. Shrader, Goddard Space Flight Center, NASA, report: “We obtained a spectrum (range 440–800 nm, resolution 1.15 nm) of the proposed optical counterpart of XTE J1859+226 (IAUC 7276) on Oct. 13.1 UT with the Steward Observatory 2.3-m telescope (+ CCD Spectropolarimeter) on Kitt Peak. The spectrum exhibits weak emission lines arising from the Balmer series of hydrogen, He II 468.6-nm, and the C III/N III blend at 464–465 nm. The Hα profile consists of a broad absorption line (FWZI = 3900 km/s) and a more-narrow emission core (FWHM = 685 km/s, equivalent width = 0.55 nm). The equivalent width of the interstellar absorption line at 578 nm is 0.11(4) nm, implying that E(B − V) = 0.8(4) mag, based on the calibration of Herbig (1975, Ap.J. 196, 129). The spectrum is typical of the optical spectra of low-mass x-ray binaries and x-ray transients and confirms the identification of the proposed optical counterpart with XTE J1859+226. Further observations are encouraged.”

A. J. Norton, C. A. Haswell, and S. Chaty, Open University; R. I. Hynes, Southampton University; A. Henden, U.S. Naval Observatory, Flagstaff; R. van Boekel, Groningen; P. M. Garnavich and K. Z. Stanek, Harvard-Smithsonian Center for Astrophysics; and T. R. Greve and J. Fynbo, Copenhagen University, report the following photometry of the proposed optical counterpart to the x-ray transient XTE J1859+226 (IAUC 7274, 7276, 7278): Oct. 12.20 UT (Fred L. Whipple Observatory 1.2-m telescope + Kitt Peak Harris filter set), U = 15.00, B = 15.75, V = 15.50, R = 15.15. Oct. 13.85 (1.0-m Jacobus Kapteyn Telescope, La Palma), Kitt Peak B = 16.03, Johnson V = 15.47, Gunn R = 15.04 (uncertainties < 0.1 mag). Oct. 13.90 (2.5-m Nordic Optical Telescope, La Palma), V = 15.31, R = 15.02, I = 14.88 (uncertainties ∼ 0.1 mag). Oct. 14.18 (USNO Flagstaff Station 1.0-m telescope), V = 15.527, B − V = +0.337, U − B = −0.647, V − R = +0.272, R − I = +0.297 (uncertainties ∼ 0.015 mag). Astrometry from the USNO observations gives α = 18^{h}58^{m}41^{s}.51, δ = +22^{o}39′30′′.2 (equinox J2000.0, referenced to USNO-A2.0).

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude estimates by A. Pearce, Nedlands, W. Australia

(0.41-m reflector): July 23.85 UT, 12.4; Aug. 21.83, 12.5; Sept. 5.87, 12.5; 15.85, 12.5; Oct. 1.84, 12.5; 10.84, 12.6.

1999 October 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green